OpenStudy (anonymous):

What do I do when I have to find derivative but I am given dx / dx (v/u) instead of (u/v) ?

OpenStudy (anonymous):

see pic of problem

OpenStudy (anonymous):

I know d(v/u) = u'v - v' u / (v^2)

OpenStudy (phi):

rename the variables in your rule: d(v/u) = u'v - v' u / (v^2) to d(u/v) = v' u - u' v/(u^2)

OpenStudy (phi):

now use the "new" rule

OpenStudy (anonymous):

ah okay. thanks!